find the distance between the parallel planes 3x-4y+6z+7=0 and 6x-8y+12z+1=0

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Student
 
First we make direction ratio same.
Hence
Planes are 
P1 : 3x -4y +6z +7 = 0
P2: 3x -4y +6z+ (1/2) = 0
 
Hence distance between parallel planes
[7 -(1/2)]/\sqrt(3² +4² + 6²)
= 13/2\sqrt61
 
Regards
Arun (askIITians forum expert)